An ideal gas is placed in a 4L container at a temperature of 300K and a pressure of 6

atmospheres. The pressure is held constant while the volume of the gas is halved. What is the

new temperature of the gas?

A. 1200K

B. 600K

C. 300K

D. 150K

E. 75K


1
Expert's answer
2022-06-10T08:28:25-0400

Given:

"V_1=4\:\rm L"

"V_2=2\:\rm L"

"T_1=300\:\rm K"


At the constant pressure the ratio of volume to temperature is constant, so

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

Thus, the final temperature

"T_2=T_1\frac{V_2}{V_1}=300*\frac{2}{4}=150\:\rm K"

Answer: D. 150K